The bearing of point A from B is 138° and the bearing of A from C is 228°. Given that the length AB=8.4cm and AC=9.8cm , calculate the bearing of B from C
Answer by Alan3354(69443)

The bearing of point A from B is 138° and the bearing of A from C is 228°. Given that the length AB=8.4cm and AC=9.8cm, calculate the bearing of B from C
----------
Angle at A = 228 - 138 = 90 degs, a right triangle
tan(C) = 8.4-9.8 = 6/7
Angle C = ~ 40.6 degs
Bearing from B to C = (228 - 180) + 40.6 = 88.6 degs
===========================